[Brahms-l] Shift report 20050226 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sat Feb 26 2005 - 16:04:21 EST
Shifter leader was SB
Shifters were 


1. There was beam from previous shift. Physics run for an hour and 
 was dumped earlier than schduled for a 2 hour machine
development work to improve beam life time. we had beam since 1:45 pm.
2. FS@6 deg. and @B861
>From previous shift run13494  175k
Day shift:
  run13498   202k
  run13499   225k
3. No detector problems so far

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Feb 26 16:04:55 2005

This archive was generated by hypermail 2.1.8 : Sat Feb 26 2005 - 16:05:09 EST